# Towards Lightweight Super-Resolution with Dual Regression Learning

### v4

@@ 1 Introduction
Deep neural networks (DNNs) have been the workhorse of many real-world applications, including image classification [1, 2] and image restoration [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Recently, image super-resolution (SR) has become an important task that aims to learn a non-linear mapping to reconstruct high-resolution (HR) images from low-resolution (LR) images. Nevertheless, the SR problem is typically an ill-posed problem and it is non-trivial to learn an effective SR model due to several underlying challenges.
First, the space of possible SR mapping functions can be extremely large since there exist many HR images that can be super-resolved from the same LR image [15]. In practice, most methods directly optimize the reconstruction loss (e.g., MAE or MSE) in HR domain and often easily obtain very blurry results with insufficient high-frequency information [16]. In other words, these undesired blurry solutions may take up the majority of the space of SR mapping to be learned and make the whole learning space extremely large. As a result, it is non-trivial to find a promising one from such a large space. To alleviate this issue, existing methods seek to increase the model capacity (e.g., EDSR [17] and RCAN [18]) and minimize the reconstruction error between the super-resolved images and the ground-truth HR images. However, these methods still suffer from such a large space of possible SR mapping functions (See analysis in Section 3.1) and often yield limited performance. Thus, how to reduce the possible space of the mapping functions to boost the training of SR models becomes an important problem.
@@ 2.1 Image Super-resolution
Existing SR methods mainly include interpolation-based approaches [24, 25, 26, 27] and reconstruction-based methods [28, 29, 18, 30, 31]. Interpolation-based methods may oversimplify the SR problem and usually generate blurry images [32, 33]. The reconstruction-based methods [34, 35, 36] reconstruct the HR images from LR images. Following such methods, many CNN-based methods [37, 38, 39, 40, 41, 42, 43] were developed to learn a reconstruction mapping.
@@ 2.2 Lightweight Model Techniques
Lightweight models have gained great attention in recent years [47, 48]. One can obtain lightweight models by directly designing efficient architectures or distilling knowledge from other models. Hui et al.[20] propose an information distillation block to extract the local long and short-path features for lightweight SR networks. Zhang et al.[49] propose a re-parameterizable building block for efficient SR. However, these models often incur a dramatic performance gap compared to state-of-the-art SR methods [21, 22]. Besides these methods, one can enhance the lightweight SR performance using knowledge distillation technique [50, 51, 52]. Gao et al.[50] use a lightweight student SR model to learn the knowledge from the deeper teacher SR network. Lee et al.[51] propose a distillation framework that leverages HR images as privileged information to boost the training of the student network. Wang et al.[53] explore the sparsity in image SR and predict a pixel-level redundancy mask to improve the inference efficiency of SR networks. Yu et al.[51] propose a hyperparameter optimization method to search for an efficient architecture for super-resolution from scratch. Yu et al.   [54] shares a similar idea with us to search/find the optimal number of channels. However, both methods define a very limited search space where all the blocks/cells share the same number of channels, ignoring the differences of redundancy among different layers. Thus, they cannot identify the layer-wise redundancy of channels and may still result in suboptimal model performance. Zhan et al.[55] further combine neural architecture search with layer-wise pruning strategy to obtain efficient SR models.
@@ 3.1 Dual Regression Learning for Super-Resolution
Figure 3: The dual regression based channel pruning method. We evaluate the importance of channels by computing both the feature reconstruction loss ${\mathcal{L}}_{\rm M}$ and the dual regression loss ${\mathcal{L}}_{\rm DR}$. Here, ${\bf X}^{(l+1)}$ and $\widehat{{\bf X}}^{(l+1)}$ denote the output features of the $l$-th layer in the original model and the pruned model, respectively. $c_{l}$ and $\hat{c}_{l}$ denote the channel number of the $l$-th layer in the original model and the pruned model.
@@ 3.2 Dual Regression Compression
Figure 4: Overview of most channel number search strategies and our importance-aware search strategy. (a) Some previous works [86, 87] assume that different channel configurations should be treated individually. For two candidate numbers of channels $k_{1}$ and $k_{2}$ ($k_{1}<k_{2}$), the selected $k_{2}$ channels are independent of the $k_{1}$ channels. (b) Some other works [88, 89, 65] use the weight-sharing strategy to reduce the search cost. For two candidate numbers of channels $k_{1}$ and $k_{2}$ ($k_{1}<k_{2}$), the selected $k_{2}$ channels contain all the $k_{1}$ channels. The weights of these overlapped channels are shared across different sets of channels during searching. (c) Based on the weight-sharing strategy, we further propose an importance-aware search strategy to search for a promising/suitable channel configuration to recognize and reduce layer-wise redundancy. For each candidate number of channels $k$, we select top-$k$ important channels and ignore the rest of the redundant channels simultaneously.
@@ Acknowledgments
This work was partially supported by Ministry of Science and Technology Foundation Project (2020AAA0106900), National Natural Science Foundation of China (62072190, 62072186), Key Realm R&D Program of Guangzhou (202007030007), Fundamental Research Funds for the Central Universities (D2191240), Program for Guangdong Introducing Innovative and Entrepreneurial Teams (2017ZT07X183), Key-Area Research and Development Program of Guangdong Province (2019B010155002), Guangdong Basic and Applied Basic Research Foundation (2019B1515130001), Tencent AI Lab Rhino-Bird Focused Research Program (JR201902), Guangzhou Science and Technology Planning Project (201904010197), Opening Project of Guangdong Key Laboratory of Big Data Analysis and Processing.

### v5

@@ 1 Introduction
Deep neural networks (DNNs) have been the workhorse of many real-world applications, including image classification [1, 2] and image restoration [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Recently, image super-resolution (SR) has become an important task that aims to learn a non-linear mapping to reconstruct high-resolution (HR) images from low-resolution (LR) images. Nevertheless, the SR problem is typically an ill-posed problem and it is non-trivial to learn an effective SR model due to several underlying challenges.
First, the space of possible SR mapping functions can be extremely large since there exist many HR images that can be super-resolved from the same LR image [19]. In practice, most methods directly optimize the reconstruction loss (e.g., MAE or MSE) in HR domain and often easily obtain very blurry results with insufficient high-frequency information [20]. In other words, these undesired blurry solutions may take up the majority of the space of SR mapping to be learned and make the whole learning space extremely large. As a result, it is non-trivial to find a promising one from such a large space. To alleviate this issue, existing methods seek to increase the model capacity (e.g., EDSR [21], RCAN [22] and ESSAformer [23]) and minimize the reconstruction error between the super-resolved images and the ground-truth HR images. However, these methods still suffer from such a large space of possible SR mapping functions (See analysis in Section 3.1) and often yield limited performance. Thus, how to reduce the possible space of the mapping functions to boost the training of SR models becomes an important problem.
@@ 2.1 Image Super-resolution
Existing SR methods mainly include interpolation-based approaches [29, 30, 31, 32] and reconstruction-based methods [33, 34, 22, 35, 36]. Interpolation-based methods may oversimplify the SR problem and usually generate blurry images [37, 38]. The reconstruction-based methods [39, 40, 41] reconstruct the HR images from LR images. Following such methods, many CNN-based methods [42, 43, 44, 45, 46, 47, 48, 49] were developed to learn a reconstruction mapping.
@@ 2.2 Lightweight Model Techniques
Lightweight models have gained great attention in recent years [53, 54, 55]. One can obtain lightweight models by directly designing efficient architectures or distilling knowledge from other models. Hui et al.[25] propose an information distillation block to extract the local long and short-path features for lightweight SR networks. Zhang et al.[56] propose a re-parameterizable building block for efficient SR. However, these models often incur a dramatic performance gap compared to state-of-the-art SR methods [26, 27]. Besides these methods, one can enhance the lightweight SR performance using knowledge distillation technique [57, 58, 59]. Gao et al.[57] use a lightweight student SR model to learn the knowledge from the deeper teacher SR network. Lee et al.[58] propose a distillation framework that leverages HR images as privileged information to boost the training of the student network. Wang et al.[60] explore the sparsity in image SR and predict a pixel-level redundancy mask to improve the inference efficiency of SR networks. Yu et al.[58] propose a hyperparameter optimization method to search for an efficient architecture for super-resolution from scratch. Yu et al.   [61] shares a similar idea with us to search/find the optimal number of channels. However, both methods define a very limited search space where all the blocks/cells share the same number of channels, ignoring the differences of redundancy among different layers. Thus, they cannot identify the layer-wise redundancy of channels and may still result in suboptimal model performance. Zhan et al.[62] further combine neural architecture search with layer-wise pruning strategy to obtain efficient SR models.
@@ 3.1 Dual Regression Learning for Super-Resolution
Figure 3: The dual regression based channel pruning method. We evaluate the importance of channels by computing both the feature reconstruction loss ${\mathcal{L}}_{\rm M}$ and the dual regression loss ${\mathcal{L}}_{\rm DR}$. Here, ${\bf X}^{(l+1)}$ and $\widehat{{\bf X}}^{(l+1)}$ denote the output features of the $l$-th layer in the original model and the pruned model, respectively. $c_{l}$ and $\hat{c}_{l}$ denote the channel number of the $l$-th layer in the original model and the pruned model. The red dashed box represents our proposed dual regression loss ${\mathcal{L}}_{\rm DR}$.
@@ 3.2 Dual Regression Compression
Figure 4: Overview of most channel number search strategies and our importance-aware search strategy. (a) Some previous works [93, 94] assume that different channel configurations should be treated individually. For two candidate numbers of channels $k_{1}$ and $k_{2}$ ($k_{1}<k_{2}$), the selected $k_{2}$ channels are independent of the $k_{1}$ channels. (b) Some other works [95, 96, 72] use the weight-sharing strategy to reduce the search cost. For two candidate numbers of channels $k_{1}$ and $k_{2}$ ($k_{1}<k_{2}$), the selected $k_{2}$ channels contain all the $k_{1}$ channels. The weights of these overlapped channels are shared across different sets of channels during searching. (c) Based on the weight-sharing strategy, we further propose an importance-aware search strategy to search for a promising/suitable channel configuration to recognize and reduce layer-wise redundancy. For each candidate number of channels $k$, we select top-$k$ important channels and ignore the rest of the redundant channels simultaneously. Note that we keep the position of the selected $k$ channels on the original model unchanged, avoiding the effect of the ranking operation on the output features.
@@ Acknowledgments
This work was partially supported by National Natural Science Foundation of China (No. 62072190, 62376099, 62072186), Key-Area Research and Development Program of Guangdong Province 2018B010107001, Program for Guangdong Introducing Innovative and Entrepreneurial Teams 2017ZT07X183, Natural Science Foundation of Guangdong Province (Grant No. 2024A1515010989).